# Position: We need responsible, application-driven (RAD) AI research

### v1

@@ 3 A Three-staged Approach to RAD-AI
Figure 1: A three-staged framework for RAD-AI, and visualisation of the dimensions and alignments that break from AI’s black box.
@@ 3.1 Laying the Foundation
Build an Integrated, Transdisciplinary Team. An interdisciplinary – even transdisciplinary (i.e. the inclusion of non-traditional academic knowledge holders) – team from the start makes accomplishing RAD-AI more feasible. By drawing on the expertise of individuals from diverse disciplines, world views, and lived experiences, the team can harness a broader range of skills, perspectives, and creativity and break out of the problem framing box. Recent calls in the AI community have focused on integrating AI with the humanities and social sciences in order to develop greater legitimacy, credibility, meaningfulness, and capability (Chades et al., 2025). Historically, incorporating domain knowledge was crucial in the success of expert systems, one of the pioneering methods of ADML systems (Confalonieri et al., 2021). Domain knowledge was the key that allowed ADML “to reason, draw new conclusions, and to generate explanations” (Confalonieri et al., 2021), legitimising its value. Its importance and success reiterate the centrality a plurality of domain expertise plays in methods development and points at the need to continue innovating on how domain knowledge is incorporated.
To advance responsible, practical black-box models, Mahinpei et al. (2021) highlights the need for better collaboration with experts to identify new, relevant concepts to focus on in intermediary quality assessments and ensure ultimate model outputs are interpretable to end users. Co-creating models with applied, transdisciplinary experts ensures practitioner-interpretable, actionable, and pragmatic outputs. Incorporating checkpoints with transdisciplinary experts also helps combat the oversimplification (i.e. problem framing box) that often occurs in AI.
Build People-Centred Studies. This step is targeted at developing transparency, trust, and interpretability. The benefits of inclusive, participatory, and people-centred research are well-recognised (Adams et al., 2024; Bednarek et al., 2025; London et al., 2020; McGrath et al., 2024; UNESCO, 2024). Two popular styles of people-centered research are human-centred design and community-based participatory research. The former is an iterative approach to aligning human desires with technologically feasible solutions, while the latter revolves around intentional, long-term relationship-building between researchers and the community (Chen et al., 2020). Both uphold values of co-creation, engagement with relevant stakeholders, and are adaptable and allow for two-way knowledge exchange (Chen et al., 2020). Depending on context, one style or a combination of the two may be more appropriate (Chen et al., 2020).
@@
Immense value to the research process comes from allowing the community “to constructively challenge, co-create, or innovate” (Stilgoe & Cohen, 2021) in the research itself. Early engagement improves the explainability of models, since “there is a gap between explainability and the goal of transparency, since explanations primarily serve ‘internal’ [computer science] stakeholders rather than ‘external’ [community] ones.” (Confalonieri et al., 2021). Further, AI’s growing ability to autonomously reach conclusions from data and provide reasonable theoretical justifications for them (Novy-Marx & Velikov, 2025) makes human oversight imperative to ensuring ethical considerations and preserving complexity of real-world systems. Hallucinations also pose a risk in these applied contexts, highlighting an ever increasing need for responsible methods development (e.g. since even small tweaks in prompting can produce different narratives and theoretical rationales from AI) alongside improvements in large language models to be more context-aware and evidence-based (Novy-Marx & Velikov, 2025).
Finally, achieving this step requires continuous, adaptive participatory engagement that evolves over time, ensuring that expectations and relationships remain resilient and aligned with the project’s goals (London et al., 2020). London et al. argue that “what makes for good community engagement is not simply the extent but the fit or alignment between the intended approach and the various contexts shaping the research projects”, which can include “the scale and scope of the …[issue], the capacities and resources of the researchers and community leaders, and the influences of the sociopolitical environment.”
@@ 3.2 Navigating Complexity
Co-Identify Application-Specific Legal and Ethical Considerations. RAD-AI research reframes research projects in a way cognisant of the context and nuances of legal and ethical condsiderations in the specific application and sector. Co-identifying these considerations involves asking ‘what is the history, context, and underlying (system-level) drivers or constrainers of the problem?’ and engaging with pre-existing law, policy, ethical codes and commitments, and public discourse. A pre-requisite for developing and assessing RAD-AI is having established principles (e.g. in the work of Seneque et al. (2024) that advances methods for aligning large language model outputs with safety and accuracy as outlined in relevant organisational AI principles). If ethical principles do not exist, these should be established here. This step also focuses on interrogating the intersection of legal and ethical considerations from different domains, as well as recognising different ethical priorities and interpretations across public, private, and non-profit sectors. Some sectors may emphasise or omit certain ethical concerns; for example, NGO and public sector documents are more likely to engage in participatory processes (Schiff et al., 2021).
@@
Examine Tacit Assumptions and Unmeasured Factors. Every study is built on tacit assumptions and unmeasured factors that influence the outcomes. Studies have been shown to reflect “the point of view that the modeler decides to capture” (Confalonieri et al., 2021), which can shape the research in ways that may not be immediately apparent. By critically examining the tacit assumptions and unmeasured factors that are brought to the study, research can build transparency (Sturmberg & Mercuri, 2024) and reflect a more comprehensive understanding of the context. For example, while a technological determinist may assume AI is both inevitable and desirable, this may not always be the case. In fact, Powles (2018) and the AAAS decision tree (n.d.) prompt the practitioner to reflect on whether AI is the right technological solution before the research begins. In this phase, researchers should ask: what am I assuming about the context, in terms of the way I am framing it and the dimensions I am capturing and omitting? What is being left unmeasured or unaddressed?
@@
Identify Application-Specific Metrics for Success. Once working in a RAD-AI setting, the need for tailored benchmarks and metrics for success in addition to standard benchmarks is paramount for evaluating efficacy in a complex, localised context (Rolnick et al., 2024; Seneque et al., 2024). Application-specific metrics for success are the relevant tests and benchmarks that can be used to appropriately assess the usability and validity of the model outputs for the intended context (Loyola-Gonzalez, 2019; Rolnick et al., 2024; Shen et al., 2023). This could include interpretable, intermediary model quality assessments and interpretable model outputs for end users (Mahinpei et al., 2021). Alternatively, Loyola-Gonzalez suggests using the Delphi method as “a statistical method for validating the suitability of an applied machine learning model by using the opinions of several experts in the application area” (2019). Other fields, such as finance, are already calling for changes in their evaluation standards in light of AI-enabled advancements (Novy-Marx & Velikov, 2025). Establishing context-specific benchmarks increases the chances that AI models are aligned with the needs of the specific application.
@@ 3.3 Testing and Sustaining Efficacy
Use Staged Testbeds. Staged, digital testbeds provide flexible, adaptable environments for testing AI systems over time, allowing researchers to identify potential hazards, and respond to changes and evolving requirements (Tzachor et al., 2022). These testbeds are an essential tool for ensuring that AI models are scalable, sensible, and responsible (Sumpter, 2024). As UNESCO (2024) notes, “testbeds should be part of the ‘agile’ regulatory toolbox available to policymakers to accommodate the fast pace of technological innovation”. By using digital testbeds, researchers can iteratively refine their models, testing them against both technical and ethical standards while evaluating their robustness in real-world contexts to ensure the results make sense in a complex world with both stated and implicit contexts. Importantly, this stage should also consider how the model and underlying training data might change with time and space. How scalable is this methodology? How are the ethics considerations, unmeasured factors, and tacit assumptions identified previously manifesting at scale? Will this research age well, and what will be its legacy in our grandchildren’s generation?
Prepare to Pass the Baton. Drawing from the leadership field, thought-leaders in the AI community have an opportunity to critically reflect on the current field and shape the legacy it will leave for the next generation. Leadership has the ability to “create, uphold, disrupt and recreate systems” (Taylor & Manning‐Ouellette, 2024); the values and ways of thinking they pass on are the ‘baton’ that future leaders start with. The question is, what is the ‘baton’ that is being passed on, and how well is the next generation equipped to carry on the legacy?
When we consider values, do we actively engage with the concerns shaping public acceptance of AI? Are we maintaining public trust in science by conducting our work with intellectual humility - the recognition of the limits of one’s knowledge (Koetke et al., 2024)? Intellectual humility manifests in many ways including acknowledging individual limitations and gaps in knowledge, actively listening to others - including the public - and reevaluating assumptions as new information emerges (Koetke et al., 2024). Testbeds provide a dual benefit here as they are a proving ground of the ‘baton’ while also offering a place to assess, refine, and iterate on what is being passed forward. They help evalutate how the next generation has been prepared, identifying both strengths and areas of improvement. Importantly, “by centering systems thinking,…leaders can pass a baton of innovative and equitable systems to the next generation” (Taylor & Manning‐Ouellette, 2024).
@@ 4 Alternative Views
Counter-positions to RAD-AI are: (1) research in artificial general intelligence (AGI), not contextual AI, is needed; (2) a checklist, not adaptive, participatory processes, is the real necessity; and (3) ‘open’ RAI is sufficient for community participation, inclusion and build trust.
@@
AGI will address our needs. While RAD-AI will allow for engaged, contextually-embedded research, it is not generalisable in the way AGI and its precursors aspire to be. This paper refers to AGI in its broadest sense, defined in the 2025 AI Safety Report as “Potential future AI that equals or surpasses human performance on all or almost all cognitive tasks” (Bengio et al., 2025, p. 218). While this scale of intelligence is yet to be achieved, even today’s most advanced AI models have concerning established and emerging risks that the most advanced mitigation techniques have not been able to overcome (Bengio et al., 2025). In particular, the gargantuan scale of the AGI learning process is near impossible to understand and obscures failure points, as learning gives rise to emergent, intermediary, and complex patterns from raw inputs. AGI also develops unforeseen capabilities (Thieme et al., 2023). This method, while powerful, often leads to a loss of explainability, making it difficult to understand how AGI systems arrive at their conclusions. Additionally, the reuse of a few foundational models in applications homogenises any defects of the original models and embeds the defects in any downstream models (Thieme et al., 2023). Rich Sutton’s “The Bitter Lesson” argues that approaches leveraging massive computation, rather than human knowledge or domain expertise, have historically yielded the most effective AI systems (2019). This framing tends to sideline critical contextual considerations of RAD-AI, reinforcing a technological deterministic viewpoint that AI can, and should, develop independently of human-driven constraints, including social, ethical, and domain-specific factors. RAD-AI, through its three-staged approach, dynamically aims to minimise such concerns in AI development.
@@
It’s time for an RAI checklist. At the other end of the spectrum, the more operationally-minded may propose implementing RAD-AI through a prescribed RAI checklist or RAI metrics (similar to (OECD.AI, n.d.)). While it may be attractive to offer ’one-and-done’ checklists, metrics, assessments, and tools, the danger is in execution: a checklist by its nature oversimplifies and reduces the nuances of context. Contrastingly, the RAD-AI framework, as well as the AAAS decision tree (n.d.) have an important distinction from a checklist - they treat the process of engaging with RAI as a continuing, dynamic conversation. The complexities of AI, including its influence across various dimensions such as societal issues, ethics, and domain-specific challenges, make the pursuit of a single defining metric impractical. As fluid as public discourse and societal concerns are, so too must be the formats of engaging with RAD-AI.
@@
‘Open’ AI is all you need. Following early efforts by Sonnenburg et al. (2007), open source has gained strong support in the research community to democratise AI innovations through platforms like GitHub. Open Source allows for transparency, reusability, and collaboration around innovations in AI, allowing methodological researchers and applied scientists to easily access, extend, and apply methods to new problems. While it may seem sufficient for RAI practices, ‘open’ AI “often lack[s] precision”, omits “scrutiny of substantial industry concentration”, and “often incorrectly appl[ies] understandings of ‘open’ imported from free and open-source software to AI systems” (Widder et al., 2024). These challenges conflict with our proposed three-staged approach that targets RAD-AI practices. Though open-source has its place, it cannot fully address RAD-AI needs.

### v2

@@ 3 A Three-staged Approach to RAD-AI
Figure 1: A three-staged framework for RAD-AI and visualisation of the dimensions and alignments that break from AI’s black box.
@@ 3.1 Laying the Foundation
Build an Integrated, Transdisciplinary Team. An interdisciplinary – even transdisciplinary (i.e. the inclusion of non-traditional academic knowledge holders) – team from the start makes accomplishing RAD-AI more feasible. By drawing on the expertise of individuals from diverse disciplines, world views, and lived experiences, the team can harness a broader range of skills, perspectives, and creativity and break out of the problem framing box. Recent calls in the AI community have focused on integrating AI with the humanities and social sciences in order to develop greater legitimacy, credibility, meaningfulness, and capability (Chades et al., 2025). Historically, incorporating domain knowledge was crucial in the success of expert systems, one of the pioneering methods of ADML systems (Confalonieri et al., 2021). Domain knowledge was the key that allowed ADML “to reason, draw new conclusions, and to generate explanations” (Confalonieri et al., 2021, p. 14), legitimising its value. Its importance and success reiterate the centrality a plurality of domain expertise plays in methods development and points at the need to continue innovating on how domain knowledge is incorporated.
To advance responsible, practical black-box models, Mahinpei et al. (2021) highlight the need for better collaboration with experts to identify new, relevant concepts to focus on in intermediary quality assessments and ensure ultimate model outputs are interpretable to end users. Co-creating models with applied, transdisciplinary experts ensures practitioner-interpretable, actionable, and pragmatic outputs. Incorporating checkpoints with transdisciplinary experts also helps combat the oversimplification (i.e. problem framing box) that often occurs in AI.
Build People-Centred Studies. This step is targeted at developing transparency, trust, and interpretability. The benefits of inclusive, participatory, and people-centred research are well-recognised (London et al., 2020; McGrath et al., 2024; UNESCO, 2024; Adams et al., 2024; Bednarek et al., 2025). Two popular styles of people-centred research are human-centred design and community-based participatory research. The former is an iterative approach to aligning human desires with technologically feasible solutions, while the latter revolves around intentional, long-term relationship-building between researchers and the community (Chen et al., 2020). Both uphold values of co-creation, engagement with relevant stakeholders, and are adaptable and allow for two-way knowledge exchange (Chen et al., 2020). Depending on context, one style or a combination of the two may be more appropriate (Chen et al., 2020).
@@
Immense value to the research process comes from allowing the community “to constructively challenge, co-create, or innovate” (Stilgoe & Cohen, 2021, p. 850) in the research itself. Early engagement improves the explainability of models, since “there is a gap between explainability and the goal of transparency, since explanations primarily serve ‘internal’ [computer science] stakeholders rather than ‘external’ [community] ones.” (Confalonieri et al., 2021, p. 15). Further, AI’s growing ability to autonomously reach conclusions from data and provide reasonable theoretical justifications for them (Novy-Marx & Velikov, 2025) makes human oversight imperative to ensuring ethical considerations and preserving complexity of real-world systems. Hallucinations also pose a risk in these applied contexts, highlighting an ever increasing need for responsible methods development (e.g. since even small tweaks in prompting can produce different narratives and theoretical rationales from AI) alongside improvements in large language models to be more context-aware and evidence-based (Novy-Marx & Velikov, 2025).
Finally, achieving this step requires continuous, adaptive participatory engagement that evolves over time, ensuring that expectations and relationships remain resilient and aligned with the project’s goals (London et al., 2020). London et al. (p. 1) argue that “what makes for good community engagement is not simply the extent but the fit or alignment between the intended approach and the various contexts shaping the research projects”, which can include “the scale and scope of the …[issue], the capacities and resources of the researchers and community leaders, and the influences of the sociopolitical environment”.
@@ 3.2 Navigating Complexity
Co-Identify Application-Specific Legal and Ethical Considerations. RAD-AI research reframes research projects in a way cognisant of the context and nuances of legal and ethical considerations in the specific application and sector. Co-identifying these considerations involves asking ‘what is the history, context, and underlying (system-level) drivers or constrainers of the problem?’ and engaging with pre-existing law, policy, ethical codes and commitments, and public discourse. A pre-requisite for developing and assessing RAD-AI is having established principles (e.g. in the work of Seneque et al. (2024) that advances methods for aligning large language model outputs with safety and accuracy as outlined in relevant organisational AI principles). If ethical principles do not exist, these should be established here. This step also focuses on interrogating the intersection of legal and ethical considerations from different domains, as well as recognising different ethical priorities and interpretations across public, private, and non-profit sectors. Some sectors may emphasise or omit certain ethical concerns; for example, NGO and public sector documents are more likely to engage in participatory processes (Schiff et al., 2021).
@@
Examine Tacit Assumptions and Unmeasured Factors. Every study is built on tacit assumptions and unmeasured factors that influence the outcomes. Studies have been shown to reflect “the point of view that the modeler decides to capture” (Confalonieri et al., 2021, p. 14), which can shape the research in ways that may not be immediately apparent. By critically examining the tacit assumptions and unmeasured factors that are brought to the study, research can build transparency (Sturmberg & Mercuri, 2024) and reflect a more comprehensive understanding of the context. For example, while a technological determinist may assume AI is both inevitable and desirable, this may not always be the case. In fact, Powles (2018) and the AAAS decision tree (n.d.) prompt the practitioner to reflect on whether AI is the right technological solution before the research begins. In this phase, researchers should ask ‘what am I assuming about the context, in terms of the way I am framing it and the dimensions I am capturing and omitting? What is being left unmeasured or unaddressed?’
@@
Identify Application-Specific Metrics for Success. Once working in a RAD-AI setting, the need for tailored benchmarks and metrics for success in addition to standard benchmarks is paramount for evaluating efficacy in a complex, localised context (Rolnick et al., 2024; Seneque et al., 2024). Application-specific metrics for success are the relevant tests and benchmarks that can be used to appropriately assess the usability and validity of the model outputs for the intended context (Loyola-Gonzalez, 2019; Shen et al., 2023; Rolnick et al., 2024). This could include interpretable, intermediary model quality assessments and interpretable model outputs for end users (Mahinpei et al., 2021). Alternatively, Loyola-Gonzalez suggests using the Delphi method as “a statistical method for validating the suitability of an applied machine learning model by using the opinions of several experts in the application area” (2019, p. 154108). Other fields, such as finance, are already calling for changes in their evaluation standards in light of AI-enabled advancements (Novy-Marx & Velikov, 2025). Establishing context-specific benchmarks increases the chances that AI models are aligned with the needs of the specific application.
@@ 3.3 Testing and Sustaining Efficacy
Use Staged Testbeds. Staged, digital testbeds provide flexible, adaptable environments for testing AI systems over time, allowing researchers to identify potential hazards, and respond to changes and evolving requirements (Tzachor et al., 2022). These testbeds are an essential tool for ensuring that AI models are scalable, sensible, and responsible (Sumpter, 2024). As UNESCO (2024, p. 45) notes, “testbeds should be part of the ‘agile’ regulatory toolbox available to policymakers to accommodate the fast pace of technological innovation.” By using digital testbeds, researchers can iteratively refine their models, testing them against both technical and ethical standards while evaluating their robustness in real-world contexts to ensure the results make sense in a complex world with both stated and implicit contexts. Importantly, this stage should also consider how the model and underlying training data might change with time and space. It asks ‘how scalable is this methodology? How are the ethics considerations, unmeasured factors, and tacit assumptions identified previously manifesting at scale? Will this research age well, and what will be its legacy in our grandchildren’s generation?’
Prepare to Pass the Baton. Drawing from the leadership field, thought-leaders in the AI community have an opportunity to critically reflect on the current field and shape the legacy it will leave for the next generation. Leadership has the ability to “create, uphold, disrupt and recreate systems” (Taylor & Manning‐Ouellette, 2024, p. 50); the values and ways of thinking leaders pass on are the ‘baton’ that future leaders start with. The question is, what is the ‘baton’ that is being passed on, and how well is the next generation equipped to carry on the legacy?
When we consider values, do we actively engage with the concerns shaping public acceptance of AI? Are we maintaining public trust in science by conducting our work with intellectual humility - the recognition of the limits of one’s knowledge (Koetke et al., 2025)? Intellectual humility manifests in many ways including acknowledging individual limitations and gaps in knowledge, actively listening to others - including the public - and reevaluating assumptions as new information emerges (Koetke et al., 2025). Testbeds provide a dual benefit here as they are a proving ground of the ‘baton’ while also offering a place to assess, refine, and iterate on what is being passed forward. They help evalutate how the next generation has been prepared, identifying both strengths and areas of improvement. Importantly, “by centering systems thinking,…leaders can pass a baton of innovative and equitable systems to the next generation” (Taylor & Manning‐Ouellette, 2024, p. 2).
@@ 4 Alternative Views
Counter-positions to RAD-AI are: (1) research in artificial general intelligence (AGI), not contextual AI, is needed; (2) a checklist, not adaptive, participatory processes, is the real necessity; and (3) ‘open’ RAI is sufficient for community participation, inclusion, and building trust.
@@
AGI will address our needs. While RAD-AI will allow for engaged, contextually-embedded research, it is not generalisable in the way AGI and its precursors aspire to be. Our paper refers to AGI in its broadest sense, defined in the 2025 AI Safety Report as “Potential future AI that equals or surpasses human performance on all or almost all cognitive tasks” (Bengio et al., 2025, p. 218). While this scale of intelligence is yet to be achieved, even today’s most advanced AI models have concerning established and emerging risks that the most advanced mitigation techniques have not been able to overcome (Bengio et al., 2025). In particular, the gargantuan scale of the AGI learning process is near impossible to understand and obscures failure points, as learning gives rise to emergent, intermediary, and complex patterns from raw inputs. AGI also develops unforeseen capabilities (Thieme et al., 2023). This method, while powerful, often leads to a loss of explainability, making it difficult to understand how AGI systems arrive at their conclusions. Additionally, the reuse of a few foundational models in applications homogenises any defects of the original models and embeds the defects in any downstream models (Thieme et al., 2023). Rich Sutton’s “The Bitter Lesson” argues that approaches leveraging massive computation, rather than human knowledge or domain expertise, have historically yielded the most effective AI systems (2019). This framing tends to sideline critical contextual considerations of RAD-AI, reinforcing a technological deterministic viewpoint that AI can, and should, develop independently of human-driven constraints, including social, ethical, and domain-specific factors. RAD-AI, through its three-staged approach, dynamically aims to minimise such concerns in AI development.
@@
It’s time for an RAI checklist. At the other end of the spectrum, the more operationally-minded may propose implementing RAD-AI through a prescribed RAI checklist or RAI metrics (similar to (OECD.AI, n.d.)). While it may be attractive to offer ‘one-and-done’ checklists, metrics, assessments, and tools, the danger is in execution: a checklist by its nature oversimplifies and reduces the nuances of context. Contrastingly, the RAD-AI framework, as well as the AAAS decision tree (n.d.) have an important distinction from a checklist - they treat the process of engaging with RAI as a continuing, dynamic conversation. The complexities of AI, including its influence across various dimensions such as societal issues, ethics, and domain-specific challenges, make the pursuit of a single defining metric impractical. As fluid as public discourse and societal concerns are, so too must be the formats of engaging with RAD-AI.
@@
‘Open’ AI is all you need. Following early efforts by Sonnenburg et al. (2007), open source has gained strong support in the research community to democratise AI innovations through platforms like GitHub. Open Source allows for transparency, reusability, and collaboration around innovations in AI, allowing methodological researchers and applied scientists to easily access, extend, and apply methods to new problems. While it may seem sufficient for RAI practices, ‘open’ AI “often lack[s] precision”, omits “scrutiny of substantial industry concentration”, and “often incorrectly appl[ies] understandings of ‘open’ imported from free and open-source software to AI systems” (Widder et al., 2024, p. 828). These challenges conflict with our proposed three-staged approach that targets RAD-AI practices. Though open-source has its place, it cannot fully address RAD-AI needs.
@@ 5 Conclusion
Acknowledgements
We thank the reviewers and our colleagues at CSIRO, especially from the Collaborative Intelligence (CINTEL) Future Science Platform (Cecile Paris), Data61 (Cara Stitzlein, Dan Pagendam, Andrea Powell, and Conrad Sanderson) and Agriculture and Food (Roger Lawes), for their insight and support throughout this work. SH was supported by a CSIRO CERC Postdoctoral Fellowship.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning in a way that contributes to a better shared future by ensuring AI research is adaptive and responsive to contextual needs and societal values.